slope:2/5
y-intercept:-3
Write the equation of the line given the following information.

Answer by confusedgirl81(14)

slope:2/5
y-intercept:-3
the equation is y = mx + b

m is the slope and b is the y-intercept
so we have
y = 2/5x + (-3)
or
y = 2/5x - 3
both equations mean the same thing!